In how many ways can three $8000 scholarships be given to three students chosen from a pool of 12 applicants?
Answer by Boreal(15235)

That is 12C3=12!/9!3!=12*11*10/6=1320/6=220 ways
Calculator 12 MATH PRB 3 then 3 ENTER